Write a detail note on applications of Differential Equation

1
Expert's answer
2022-03-31T08:45:25-0400

Differential equations are important as they are used in calculation of the rates of change of various systems

E.g.

1.Exponential population growth

Let P(t) be a quantity increasing with time, t and the rate of increase is proportional to P

I.e.

"\frac{dP}{dt}=kP"

Solution to this differential equation is given by

"P(t)=A e^{kt}"



2. Exponential Decay- Radioactive Material

Let M(t) be the amount of a product decreasing with time, t and decreasing rate is proportional to M

I.e.

"\frac{dM}{dt}=-kM"


Where, "\frac{dM}{dt}" is the first derivative of M

Solving this we get,

"M(t)=A e^{-kt}"

M(t) = A e- k t


3. Falling Object

An object is dropped from a height at time t = 0. If h(t) is the height of the object at time t, a(t) the acceleration and v(t) the velocity. The relationships between a, v and h are as follows

"a(t)=\frac{dv}{dt}, v(t)=\frac{dh}{dt}"



4. Newton's Law of Cooling

It is a model that describes, mathematically, the change in temperature of an object in a given environment. The law states that the rate of change (in time) of the temperature is proportional to the difference between the temperature T of the object and the temperature Te of the environment surrounding the object.


"\frac{dT}{dt}=-k(T-Te)"


Let x= T - Te hence, "\frac{dx}{dt}={dT}{dt}"

Using this, the above formula becomes

"\frac{dx}{td}=-kx"

Hence we have, "x=A e^{-kt}"


The solution to the above differential equation is given by

x = A e - k t

substitute x by T - Te

T - Te = A e - k t

Assume that at t = 0 the temperature T = To

To - Te = A e 0

which gives A = To - Te

The final expression for T(t) i given by

T(t) = Te + (To - Te)e - k t

This last expression shows how the temperature T of the object changes with time.






5. RL Circuit


Consider an RL circuit

At t = 0 the switch is closed and current passes through the circuit. Electricity laws state that the voltage across a resistor of resistance R is equal to R i and the voltage across an inductor L is given by "L\frac{di}{dt}" (I is current).